Write an equation for the line described. Give your answer in standard form. x-intercept -5, y-intercept 2. Answer is -5x + 2y=10. I am not sure why this problem is equal to 10. I know the line standard form is based on Ax+By=C. Thank you for your help.
Answer by mananth(16949)

x1 y1 x2 y2
-5 0 0 2

slope m = (y2-y1)/(x2-x1)
( 2 - 0 )/( 0 - -5 )
( 2 / 5 )
m= 2/ 5

Plug value of the slope and point ( -5 , 0 ) in
Y = m x + b
0.00 = -2 + b
b= 0 - -2
b= 2
So the equation will be
Y = 2/5 x + 2
multiply by 5
5y=2x+10
-2x+5y=10

m.ananth@hotmail.ca